Solve the equation for y. Then find the value of y for the given value of x.
3x-6y=6; x=2

Answer by abdulrasid84@gmail.com(4)

3(2)-6y=6
6-6y=6
-6y=6-6
-6y=0
o/-6=0
then the value of y is equal to zero
y=0